how do you solve this triangle 24in by 15in by 19in and get the answer 143in to the second power
Answer by stanbon(75887)

how do you solve this triangle 24in by 15in by 19in and get the answer 143in to the second power
---
Area = (1/2)base*height
---
If 143 is area, base = 15 and height = 19, you get:
Area = (1/2)15*19
Area = (1/2)*285
---
Area = 142.5
----
Rounded up you get:
Area = 143 in^2
======================
Cheers,
Stan H.
=============